What is the answer in a solution:two ninth of the people in a restaurant are adults.if there 91 adults and children,how many children are there in the restaurant..thankyou.here's my email tatabalabs@gmail.com
Answer by stanbon(75887)

What is the answer in a solution:two ninth of the people in a restaurant are adults.if there 91 adults and children,how many children are there in the restaurant.
----------
Let "a" be the number of adults.
Let "c" be the number of children
Equation:
If the number of adults is 2/9, the number of children
must be 7/9
-----
c = (7/9)91 = 7*(10 1/9) = 71 when rounded up
-------
Cheers,
Stan H.
-------------
--------